Shift report 20020106 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sun Jan 06 2002 - 08:59:28 EST

  • Next message: Flemming Videbaek: "Trigger ratios etc."

    ------------------------------------------------------------
    Date: Sun Jan  6 08:59:28 2002, Shift: 00:00-08:00
    Supervisor: sjs, with: 
    Log book#12 pages 6 - 7
    
    Shift Summary:
    ------------------------------------------------------------
    (2nd attempt after first shift report disappeared into
    byte heaven...)
    
    1.) Continuous store until ~7:20.  MCR called at 7:10 asking
    if we would be willing to dump indicating the other experiments wanted a dump.  I agreed since we had just completed the FS setting and needed to do a shift check. 
    (STAR called about 2 minutes later asking if we REALLY 
    wanted to dump....)  Currently in a "quench link interlock"
    situation...
    
    2.) Two long data runs overnight:
      6302 (MRS 12 deg 4B) 123K effective trig 3 (615K actual)
                            42K effective trig 6 (85K actual)
    
      6303  (MRS 12 deg 4A) 122K effective trig 3 (611K actual)
                             42K effective trig 6 (85K actual)
    
    3.) DVM spectra not being generated correctly during data
    runs.  However, the DVM spectrum obtained from the 
    calibration run looks OK.  (Is this related to the other
    trigger problems?)
    
    Towards end of run 6303 discovered that BRAVO would
    not read out TOFW, H1, H2, and C1, complaining that it was
    "Unable to connect to mainframes".  There is nothing 
    obvious in the data to suggest a problem so this looks
    like a communication problem of some sort...
    
    Although not noted in the log book, the Si detector electronic was reconfigured last evening to bypass the PTQ units.  For the low amplitude signals that we have for the pp runs, this seems to give better separation of the single MIP peak from the pedestal.  It is noted, however, that several channels now have bad DC offsets which should be
    fixed...
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sun Jan 06 2002 - 09:00:14 EST